May 1st, 2016
For Sales a BSS BLU-160 signal processor recently removed from a local theater in Singapore. It is working fine and updated to firmware 86.02.02. It come with two input and two output analogy cards thus with total 8 input and 8 output channels.
Picture are the units and I am selling it for S$1050 or US$750. Shipping will be additional US$100 by EMS for overseas. Additional picture can be provided. Used and is good working conditions. Only have some scratches due to rack mount before but no big deal.
This is the machine to use for JBL M2 and 7 series monitor speakers. Interested please send mail to me.
February 1st, 2016
Toshiba K170BL – only 6.5mA to 9mA (HK$100 for 4 pcs)
Toshiba J74BL – 6.4 to 9.8mA (HK$110 each pc); 9.9 to 11.2mA (HKD$75 each pc)
Toshiba K170 / J74 pair HK$150 each pair.
2SJ162 / 2SK1053 (Hitachi) : 22 pairs left (HKD$200 per pair)
2SJ200 / 2SK1529 (Toshiba): 20 pairs left (HK$180 per pair)
FDA-2A assembly DAC x 1 (older version) HK$1500 free shipping. Fully tested.
Note that price need to add shipping and paypal fee unless otherwise specified free.
December 12th, 2015
These were newly acquired in Nov 2015 from eBay.
PXIe-1062Q: for both PXI and PXIe cards with power supply.
PXIe-8105: controller for the system with upgrade to IDE 120 GB SSD from Kingspec (China) and 4 GB (2 x 2GB) DDR2 RAM. Actually I found all on hand DDR2 ram with higher speed with work on this controller. BIOS has also been updated to the latest version 1.3.3. Speed wise is ok fast now with SSD and 4GB ram under window 7 32bit.
PXI-4130: SMU with two output: +/-20V max 2A and 9V 1A. This supply can be programmed to check transistors or FETs under Labview.
The total cost is about USD$1500 but if getting a SMU with dual output will be much more expensive!
I shall learn and try to develop a program to sort the mosfet 2SK1530 / 2SJ201 soon. Sorting of the Toshiba 2SK170 / J74 is also very good as the speed is supper fast and reading is stable with no oscillation concern.
USB GPIB adaptor 82357B from Agilent has big issue to link up with other equipment on the NI controller even I follow the steps as what is advised by Aglient website. Finally I get the original NI GPIB cable Type X13 183285-02 and every things works perfectly. No wonder the price of second hand USB-GPIB cable from Agilent is half of that from NI!
Changing the 3V back up battery in the PXIe-8105 controller must need reset the reset jumper pins or else the controller cannot boot up at all.
December 12th, 2015
The is a meter used in DI water conductivity monitoring and failure symptom is no display. Quick analysis show that the Switch Mode power supply inside has no output. Thus it should be an easy fix but …
The switch mode power supply provide about 10W to the controller and thus it is stack on top of bigger pcb using some pin header. The AC main filter, fuse, current limit etc are house on the bottom bigger pcb and the whole SWPS is on the small PCB. The design is using a DIP-7 pin IC as the switch but the part number of it is masked (removed). Thus the circuit need to be traced by reverse engineering as shown on the draft. After searching on internet, the only match part is from ROHM BM2Pxx4 series IC. Thus I decided to get in the BM2P014 for trial. Other problem seen in the power supplier PCB is one 10u 400V capacitor swell on top and after removing it found capacitor value 99% gone. Thus I also order two Nichicon 105 degC grade for replacement. When I first replace the IC BM2P014, the supplier seems to output voltage but about 2 minutes, there is no more output and thus the NEW IC is spoil! So left only one ICs on hand a bit worry not to get it repaired! After looking around, I remove the capacitor at pin 5 (Vcc supplier filter) and also found no capacitance value! Thus I concluded that the failure is caused by poor filter of the Vcc supply and then kill the ICs after some time. Due to that I also change all the E-cap on the output side of the supply and measurement show that most of the capacitors are out of spec with lower capacitor value due to prolong use of the meter (non stop). I just wonder why the meter use such low grade unbranded so call 105 degC capacitors inside this industry meter! The cost to get replacement for this meter is about US$650 !!!
SMPS with DIP-7 IC and two 10u 400V capacitors removed:
Draft circuit of the SMPS:
Components ordered from Mouser:
Defective parts from SMPS:
Final repaired power supply:
Note the output of power supply is +12V 0V -12V; Ve is oV and Vf is 24V.
October 17th, 2015
This scope failed 3 years ago and the symptom was just no display after power on but some LEDs flicks to a point and stopped at “ADD” LED there! Recently I take it out and attempt to repair it.
The first suspect is the high voltage board defective and I open it and do a check on the high voltage board switching transistor and everything looks fine. Also I can get about 65k switching frequency at the switch mode power supply. After looking around, I see there is a “G Bias” pot on the high voltage board and then I turn it clockwise and see some display DOT on the screen. The scope wordings of time base and input attention can be seen also. Thus I reboot it and see a “Step 05 fail 44” error on the screen. After checking the service manual it point to the DAC on the control board and thus I look at it again. Oh I find the leaking SMD capacitors with corrosion in its surrounding components.
It can be seen that the components surrounding the three E-cap are with corrosion solder join! So in totally changed all 4 SMD e-caps on the control board as the other corner also have one more leakage e-cap.
After changed the 4 caps, I put back the board and power on the set but the failure symptom is still same step 05 fail 44!!! So I probe around the DAC chip and check the Vref 10V at pins 14 and pin 15. The voltage is really low and close to 0 volt.
So I started to check the components around the DAC U2234 and later find out that the resistor R2228 10k is open circuit. I replace it with another one smd resistor and after that I power up the set and it works again!
I burn in the set for about 15 mins and then I start to smell something from the scope and of course power off it immediately. Open up and found the line filter caps (RIFA) brands are all burnt together with two more resistors.
C1018, R1018, C1016, R1010 (parallel with RT1010; 15 ohm not on schematic) were all burnt of open circuits. Later I also find three more Rifa capacitors on the other PCB (2200pf and 10nF 250v ~ is also crack a bit on the body. I take them out and all the 2200 pf caps are with very low value of about 100pf! Thus I order them from element14 and they arrive after 3 days. For the 0.068u 250V~ cap, I use on hand 0.1uf 250V~ cap to replace them (blue color on the picture). After put back, everything works fine. I have brought this 25 years old scope back to life!
I also note that there are many same scope 2445 or 2445B with similar failure of step 05 fail 44. I hope this information can help some of the people to repair their scope.